OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the following circle. x^2 + y^2 = 16

OpenStudy (hoblos):

can you try this one? it's easy

OpenStudy (anonymous):

ok so whats my steps aqain

OpenStudy (hoblos):

if the center is (x1 , y1) and the radius is r then the equation is (x-x1)^2 + (y-y1)^2 = r^2

OpenStudy (hoblos):

now compare this equation with the equation you have to figure out the center and the radius

OpenStudy (anonymous):

so how would i do it aqain?

OpenStudy (hoblos):

the equation you have can be written as (x-0)^2 + (y-0)^2 = 4^2 do you get it know??

OpenStudy (anonymous):

so 1,1 16

OpenStudy (hoblos):

it is (0,0) 4

OpenStudy (anonymous):

thnx

OpenStudy (hoblos):

my pleasure XD hope you understood it